# Maximum-likelihood estimation of the Matérn covariance structure of isotropic spatial random fields on finite, sampled grids

### v1

@@ 4.3 Finite sampled data: full theory
For sampled data the likelihood involves the blurred spectral density $\bar{\mathcal{S}}_{\boldsymbol{\theta}}(\mathbf{k})$, and the modified periodogram of the data ${\mathcal{H}}(\mathbf{x})$,
@@
The solution $\hat{\theta}$ is found by maximization of eq. (48), requiring the vanishing of the score $\mbox{\boldmath$\nabla$}{\bar{{\mathcal{L}}}(\boldsymbol{\theta})}={\mbox{\boldmath$\bar{\gamma}$}}(\boldsymbol{\theta})$, whose elements are given by
@@
with the blurred equivalents to eq. (41) again obtained exactly via the intermediary of eq.( 50) as
@@
As [Simons & Olhede(2013)] (their eq. 138), but now following [Guillaumin et al.(2022)Guillaumin, Sykulski, Olhede, & Simons] (their eq. 36), the parameter estimation variance,
@@
requires the additional calculation of the covariance of the score without neglecting the correlation between wavenumbers,
@@
see [Guillaumin et al.(2022)Guillaumin, Sykulski, Olhede, & Simons] (their eq. 37), and compare [Simons & Olhede(2013)] (their eq. A54, which should have quoted blurred quantities), which implies that we require the covariance of the windowed periodogram $|H(\mathbf{k})|^{2}$, which under standard theory using the [Isserlis(1918)] theorem for Gaussian processes is [[]their eq. A57]Percival+93,Stein95,Simons+2013
@@ 4.4 Covariance of the estimates
The first is by direct calculation of the covariance between the windowed periodograms in eq. (57), rewritten [Walden et al.(1994)Walden, McCoy, & Percival] as
@@
where, abusing notation ever so slightly, $\mathsf{Q}$ is the discrete-Fourier transform matrix operator that transforms windowed spatial-domain observations to the spectral domain, $\mathsf{Q}^{\prime}$ and $\mathsf{Q}^{*}$$^{\prime}$ its transpose and conjugate transposes, respectively, and ${\mathcal{C}}_{\boldsymbol{\theta}}$ the Matérn autocovariance sequence at all available lags. Fig. 6 illustrates this computational procedure for a spatial observation grid with a unit-window taper. The top row shows the covariance sequence ${\mathcal{C}}_{\boldsymbol{\theta}}(\|\mathbf{x}-\mathbf{x^{\prime}}\|)$ with all two-dimensional lag-distances unwrapped along the two axes labeled $\mathbf{x}(:)$ and $\mathbf{x^{\prime}}(:)$, followed by the covariance of the periodograms as the sum of the individual (pseudo-)covariance [Neeser & Massey(1993)] terms as they appear in eq. (57). The bottom row shows the $\mathbf{k}=\mathbf{k^{\prime}}$ diagonal elements of each of the latter, wrapped as $k_{x}$ and $k_{y}$ about the zero wavenumber in the center, following the independently (via eq. 50) calculated blurred square of the spectral density, as labeled. Indeed, as per eq. (30), the main diagonal of the periodogram covariance, its variance, equals the square of the blurred spectral density, that is, $\bar{{\mathcal{S}}}^{2}\hskip-1.00006pt(\mathbf{k})=\mbox{var}\big{\{}|H(\mathbf{k})|^{2}\}$.
@@
$\langle H(\mathbf{k})H^{*}\hskip-1.00006pt(\mathbf{k^{\prime}})\rangle=\int\!\!\!\!\int D_{K}(\mathbf{k}-\mathbf{k}^{\prime\prime})D_{K}^{*}(\mathbf{k}^{\prime}-\mathbf{k}^{\prime\prime})\hskip 1.00006pt{\mathcal{S}}_{\boldsymbol{\theta}}(\mathbf{k}^{\prime\prime})\,d\mathbf{k}^{\prime\prime}\quad\mbox{where}\quad D_{K}(\mathbf{k})=\frac{1}{2\pi}\left(\frac{\Delta x\Delta y}{MN}\right)^{\frac{1}{2}}\sum_{\mathbf{x}}w(\mathbf{x})e^{-i\mathbf{k}\cdot\mathbf{x}}.$ (61)
@@
$\langle H(\mathbf{k})H(\mathbf{k^{\prime}})\rangle=\int\!\!\!\!\int D_{K}(\mathbf{k}-\mathbf{k}^{\prime\prime})D_{K}(\mathbf{k}^{\prime}+\mathbf{k}^{\prime\prime})\hskip 1.00006pt{\mathcal{S}}_{\boldsymbol{\theta}}(\mathbf{k}^{\prime\prime})\,d\mathbf{k}^{\prime\prime}.$ (64)
This procedure requires one Fast Fourier Transform per diagonal increasingly offset from the main diagonal in $(\mathbf{k},\mathbf{k^{\prime}})$ space, further multiplied with $\bar{m}_{\theta}(\mathbf{k})/\bar{\mathcal{S}}_{\boldsymbol{\theta}}(\mathbf{k})$ and $\bar{m}_{\theta}(\mathbf{k^{\prime}})/\bar{\mathcal{S}}_{\boldsymbol{\theta}}(\mathbf{k^{\prime}})$ and normalized into eq. (56). Like the procedure in eqs (58)–(60), it is exact [[, as opposed to Riemann discretization of eqs (61) and (64), as proposed by]]Guillaumin+2022.
Figure 6: Sampling and windowing effects on the covariance between periodograms: blurring and correlation illustrated for a Matérn model with parameters $\boldsymbol{\theta}_{0}=({\sigma^{2}},\nu,\rho)=(1,0.5,2)$ on a grid of $M=N=9$ pixels of size $\Delta x=\Delta y=1$ in arbitrary units. (Top row) The spatial covariance, ${\mathcal{C}}_{\boldsymbol{\theta}}(\mathbf{x},\mathbf{x^{\prime}})$, on the grid that gives rise to the periodogram covariance, $\mbox{cov}\big{\{}|H(\mathbf{k})|^{2},|H(\mathbf{k^{\prime}})|^{2}\big{\}}$, which is the sum of two (pseudo-)covariance terms. The off-diagonal terms express the correlation between periodograms at different frequencies; the diagonal terms embody their blurring. (Bottom row) The squared blurred spectral density $\bar{{\mathcal{S}}}^{2}\hskip-1.00006pt(\mathbf{k})$ is equal to the variance of the periodograms, $\mbox{var}\big{\{}|H(\mathbf{k})|^{2}\}$, as shown in the last three panels, which are identical to the diagonals of the panels above them, wrapped into a two-dimensional spectral grid.
Assembling all the various pieces, we now have three different methods to obtain the desired estimation variance in eq. (56): likelihood gradient sampling, direct calculation via the DFT matrix, and per-diagonal Fourier transformation of the spatial covariance. The first is fast and has shown good behavior in all of our tests, see Sec. 4.5. The second and third produce identical results, with the matrix implementation being memory intensive, and the per-diagonal sequential calculation still computationally heavy as it calls on $MN$ FFTs, each of which requiring $MN\log(MN)$ operations. As shown in Fig. 6, not all interaction terms are large, which may inspire future approaches utilizing Markov-Chain Monte Carlo approaches or other forms of importance sampling [Metropolis & Ulam(1949), Hastings(1970)]. Managing the interplay of the stochastic process being sampled and the characteristics of the sampling strategy is amenable to formal treatment via the significant correlation contribution as developed by [Guillaumin et al.(2017)Guillaumin, Sykulski, Olhede, Early, & Lilly, Guillaumin et al.(2022)Guillaumin, Sykulski, Olhede, & Simons].
@@ 4.5 Synthetic numerical examples
Fig. 7 is a representative illustration of the behavior of the estimator $\hat{\theta}$ of eqs (46)–(48), in each of the Matérn parameters $\sigma^{2}$}, $\nu$, and $\rho$. We conducted five hundred inversions on independently generated simulations (via circulant embedding of the spatial covariance), and studied the empirical distribution of the estimates obtained using exact blurring per eqs (52)–(53), compared to the theoretical expression of eq. (55) using the exact method of eqs (58)–(60). Invariably, the estimates are nearly unbiased, and nearly universally Gaussian distributed, as can be seen from the histograms and the quantile-quantile plots. The top row shows the smoothly estimated [Botev et al.(2010)Botev, Grotowski, & Kroese] standardized probability density functions (shaded) of the values recovered in this experiment of sample size $64\times 64$. The abscissas were truncated to within $\pm$3 of the empirical standard deviation about the origin which is the mean estimate; the percentage of the values captured by this truncation is listed in the top left of each graph. The thick gray line summarizes the empirical distributions as a Gaussian with the observed mean and variance in this set of simulations. The thick black line uses the corresponding theoretical values calculated from eq. (55). The ratio of the empirical to the theoretical standard deviation is shown listed as $s/\sigma$ for each of the parameters. The dotted black line marks the true parameter values. The bottom row shows the quantile-quantile plots of the theoretical (abscissa, horizontal) versus the empirical (ordinate, vertical) distributions. The averages of the recovered values ${\sigma^{2}},\nu$ and $\rho$ are listed at the top of the second row of graphs. The true parameter values $\sigma^{2}_{0},\nu_{0}$ and $\rho_{0}$ are listed at the bottom. We attribute the slight asymmetry of the histograms and the small bends at the edges of the quantile-quantile plot to small sample effects on minimum/maximum order statistics, and to all Matérn parameters being necessarily positive.
Figure 7: Matérn parameter maximum-likelihood estimation statistics for 500 spatial-covariance embedding simulations carried out on a $\times64$ grid, with spacings $\Delta x=\Delta y=10$ km, and true values of $\sigma^{2}_{0}=1$ km2, $\nu_{0}=2.5$, $\rho_{0}=20$ km, recovered via maximization of the exactly blurred uncorrelated likelihood (48). The estimates average to ${\sigma^{2}}=0.98\pm 0.27$ km2, $\nu=2.56\pm 0.20$, $\rho=19.64\pm 1.99$ km, quoting one observed standard deviation. The thick gray line is derived from the covariance calculated from the ensemble of simulation and estimation outcomes. The thick black line is based on the covariance exactly calculated from eq. (55) shown and calculated at the truth, shown by the dotted vertical line in the top panels.
@@
Figure 8: An example of a likelihood (48) surface for simulated data from the experiment reported in Fig. 7 that landed an estimate very close to the truth. Every panel fixes the third parameter that is not shown at its estimate. Likelihood contours are drawn and shaded in ten equal increments up until the likelihood that corresponds to the one reached by the parameter pair that is twice the empirical standard deviation away from the estimate, in the center of each panel.
Figure 9: Maximum-likelihood estimation statistics for the ensemble of 500 simulation and recovery experiments reported in Fig. 7. The mean estimate is highlighted by the gray triangle and two observed standard deviations marked gray by lines. The heavy gray ellipse is the 68% confidence region based on the ensemble. The heavy black ellipse is the predicted 68% confidence region based on the covariance predicted from eq. (55), which is a close match.
Figure 10: Comparison of the covariance predicted via eq. (55) and the covariance observed on the set of experiments reported in Figs. 7 and 9. Shown are the relevant correlation matrices between the estimators for the three Matérn parameters $\sigma^{2}$}, $\nu$, and $\rho$, highlighting the relatively strong trade-off between $\nu$ and $\rho$.
Fig. 9 summarizes the same experiment as cross plots of the recovered parameters with their empirical summaries [Aster et al.(2013)Aster, Borchers, & Thurber] as the ellipsoidal 68% confidence region (thick gray line), and showing the equivalent error ellipses from the theoretically predicted covariances (thick black line), which match the observations closely, though not perfectly. Every panel shows all pairs of parameter estimates (gray circles), with the mean marked as a gray triangle, and a cross indicating twice the standard deviation in each of the estimates. In the lower left of each graph we display the percentage of the estimates that falls within the error ellipse calculated for the observed covariance.
Fig. 10 is an alternative graphical rendition of the theoretically predicted and empirically observed parameter-estimate correlation matrix, that is, their covariance matrix normalized to unity along the diagonal. Parameter trade-offs are inherent in the Matérn description (see, for example, eq. 14). The theoretically calculated correlations are $\{{\sigma^{2}},\nu\}=-0.2837$, $\{{\sigma^{2}},\rho\}=0.8167$, and $\{\nu,\rho\}=-0.6766$, which are good predictors for those observed, as the empirical values are $\{{\sigma^{2}},\nu\}=-0.1961$, $\{{\sigma^{2}},\rho\}=0.8335$, and $\{\nu,\rho\}=-0.5678$, respectively.
@@ 4.6 Analysis of residuals
The terms ${\mathcal{S}}_{\boldsymbol{\theta}}^{-1}\hskip-1.00006pt(k)\hskip 1.00006pt|H(\mathbf{k})|^{2}$ and $\bar{\mathcal{S}}_{\boldsymbol{\theta}}^{-1}\hskip-1.00006pt(k)\hskip 1.00006pt|H(\mathbf{k})|^{2}$ that have appeared above in the expressions for the likelihoods (39), (48), and their derivatives (40) and (52) contain the ratio of the observed periodogram of the data to the (blurred) spectral density predicted under the model. Since the spectral density is a ‘scale’ parameter (as opposed to a ‘location’) this ratio has the usual interpretation as a measure of misfit. We have previously noted that if the Gaussian model fits, the expectation of this quadratic is $\langle\bar{\mathcal{S}}_{\boldsymbol{\theta}}^{-1}\hskip-1.00006pt(k)\hskip 1.00006pt|H(\mathbf{k})|^{2}\rangle=1$. Here we follow [Simons & Olhede(2013)] to maintain that twice this quantity, a quadratic form of Gaussian random variables, given that ${\mathcal{H}}(\mathbf{x})$ and thus $H(\mathbf{k})$ are assumed to be Gaussian and uncorrelated, should be a chi-squared random variable with two degrees of freedom,
$X_{\boldsymbol{\theta}}(\mathbf{k})=\bar{\mathcal{S}}_{\boldsymbol{\theta}}^{-1}\hskip-1.00006pt(k)\hskip 1.00006pt|H(\mathbf{k})|^{2}\sim\chi^{2}_{2}/2.$ (65)
@@
It is imprudent to ignore and impossible to overstate the importance of such a hypothesis test. Apart from serious numerical instability and potential run-away effects, possibly caused by improper initialization of or unrealistic constraints on the optimization procedure, maximum–likelihood inversion will always return the parameter set with maximum likelihood. But whether the most likely model is, in fact, any good, then remains to be ascertained. Establishing whether eq. (65) in fact holds can be carried out visually, by inspection of the overlay of the histograms of $2X_{\boldsymbol{\theta}}$ across all wave vectors with the probability density function $\chi^{2}_{2}$, and by making ‘quantile-quantile’ plots of the ranked values of $2X_{\boldsymbol{\theta}}$ versus the inverse cumulative density function of $\chi^{2}_{2}$ evaluated at their corresponding fractional ranks. Moreover, the two-dimensional map of $X_{\boldsymbol{\theta}}(\mathbf{k})$ should show no residual structure, and will contain information on possible wavenumber ranges or specific directions (for example, in the presence of anisotropy) in which the data might be over- or under-fit. All three such representations of model quality (histograms, quantile-quantile ‘Q-Q’ plots, and two-dimensional wavenumber map of the residuals) must be thoroughly scrutinized.
Beyond visual inspection it is desirable to design a formal test for when the hypothesis of isotropic Matérn behavior needs to be abandoned, and the veracity of the parameters recovered by likelihood maximization called into question, regardless of how narrow their uncertainty intervals (55) may be. Failing the test could be due, for example, to non-Gaussianity, or to the presence of patterns or preferred directions indicating that the data should rather be interpreted under anisotropic [[, e.g.,]]Goff+89b,Herzfeld+99,Olhede2008,Olhede+2014b extensions of the model. We save developing alternative hypotheses for future work.
@@
The smoothness and boundedness of the spectrum $\bar{\mathcal{S}}_{\boldsymbol{\theta}}(k)$, the presumed independence of $X_{\boldsymbol{\theta}}(\mathbf{k})$ between wavenumbers, and the central limit theorem should help the variable $s^{2}_{X}$ to converge to a normal variate. The central moments of the $p$th power of chi-squared variables with $m$ degrees of freedom [Davison(2003)] satisfy $\big{\langle}\!\left[\chi^{2}_{m}\right]^{p}\!\big{\rangle}=2^{p}\Gamma(p+m/2)/\Gamma(m/2)$, from which we obtain
@@
The case $p=1$ discussed previously, in reducing eq. (42) to eq. (43), is easily verified. Evaluating eq. (67) for the case $p=2$ then yields the expectation of eq. (66), our test statistic,
@@
Hence we deduce that our chosen metric converges ‘in law’ [Ferguson(1996)] to a variable distributed as:
@@
Figure 11: Residual statistics of simulations carried out on a $\times64$ grid, with spacings $\Delta x=\Delta y=10$ km, convolutional blurring, and true Matérn parameter values of $\sigma^{2}_{0}=1$ km2, $\nu_{0}=2.5$, $\rho_{0}=20$ km. Results pertaining to one of the simulations and its maximum-likelihood recovery. Distribution of the variable $X_{\boldsymbol{\theta}}(\mathbf{k})$ of eq. (65), as a histogram across all wavenumbers with the theoretical distribution superposed, as a quantile-quantile plot for the distribution in question, and as a spectral-domain map.
@@
Fig. 12 validates the distribution of the test statistic through a series of simulation experiments with the same parameters as those used for Fig. 11. Reporting on the ‘test of the test’, specifically, whether eq. (71) holds, the two panels illustrate the distribution of $s^{2}_{X}$ across 175 realizations. Sample mean and variance of the test statistic are labeled above the first panel, which displays the histogram of $s^{2}_{X}$ over the simulations. The second panel shows the linearity of the quantile-quantile plots. Our conclusion is that using $s^{2}_{X}$ as a statistic results in a useful and sensitive test on the appropriateness of the Matérn model, whatever its parameters, and irrespectively of their confidence intervals.

### v2

@@ 4.3 Finite sampled data: full theory
For sampled data, the likelihood involves $\bar{\mathcal{S}}_{\boldsymbol{\theta}}(\mathbf{k})$, the blurred spectral density (eqs 24 and 30), and the modified periodogram of the data ${\mathcal{H}}(\mathbf{x})$,
@@
The solution $\hat{\theta}$ is found by maximization of eq. (48), requiring the vanishing of the score $\mbox{$\nabla$}_{\!{}_{\boldsymbol{\theta}}}{\bar{{\mathcal{L}}}(\boldsymbol{\theta})}={\mbox{$\bar{\gamma}$}}(\boldsymbol{\theta})$, whose components are given by
@@
with the blurred equivalents to eq. (41) again obtained exactly via the intermediary of eq. (50) as
@@
As [58] (their eq. 138), but now following [28] (their eq. 36), the parameter estimation variance,
@@
where in practice we substitute $\boldsymbol{\hat{\theta}}$ for $\boldsymbol{\theta_{0}}$, to which the estimator converges in probability, requires the additional calculation of the covariance of the score without neglecting the correlation between wavenumbers,
@@
see [28] (their eq. 37), and compare [58] (their eq. A54, which should have quoted blurred quantities). Eq. (56) implies that we require the covariance of the windowed periodogram $|H(\mathbf{k})|^{2}$, which under standard theory using the [38] theorem for Gaussian processes is [[]their eq. A57]Percival+93,Stein95,Simons+2013
@@ 4.4 Covariance of the estimates
The first is by direct calculation of the covariance between the windowed periodograms in eq. (57), rewritten [65] as
@@
where, abusing notation ever so slightly, $\mathsf{Q}$ is the discrete-Fourier transform matrix operator that transforms windowed spatial-domain observations to the spectral domain, $\mathsf{Q}^{\prime}$ and $\mathsf{Q}^{*}$$^{\prime}$ its transpose and conjugate transposes, respectively, and ${\mathcal{C}}_{\boldsymbol{\theta}}$ the Matérn autocovariance sequence at all available lags. Fig. 6 illustrates this computational procedure for a spatial observation grid with a unit-window taper. The top row shows the covariance sequence ${\mathcal{C}}_{\boldsymbol{\theta}}(\|\mathbf{x}-\mathbf{x^{\prime}}\|)$ with all two-dimensional lag-distances unwrapped along the two axes labeled $\mathbf{x}(:)$ and $\mathbf{x^{\prime}}(:)$, followed by the covariance of the periodograms as the sum of the individual (pseudo-)covariance [46] terms as they appear in eq. (57). The bottom row shows the $\mathbf{k}=\mathbf{k^{\prime}}$ diagonal elements of each of the latter, wrapped as $k_{x}$ and $k_{y}$ about the zero wavenumber in the center, following the independently (via eq. 50) calculated square of the blurred spectral density, as labeled. The main diagonal of the periodogram covariance, its variance $\mbox{cov}\big\{|H(\mathbf{k})|^{2},|H(\mathbf{k})|^{2}\big\}=\mbox{var}\big\{|H(\mathbf{k})|^{2}\}$, matches the square of the blurred spectral density $\bar{{\mathcal{S}}}_{\boldsymbol{\theta}}^{2}\hskip-1.00006pt(\mathbf{k})$.
@@
$\langle H(\mathbf{k})H^{*}\hskip-1.00006pt(\mathbf{k^{\prime}})\rangle=\iint\limits_{\mathbf{k}}D_{K}(\mathbf{k}-\mathbf{k}^{\prime\prime})D_{K}^{*}(\mathbf{k}^{\prime}-\mathbf{k}^{\prime\prime})\hskip 1.00006pt{\mathcal{S}}_{\boldsymbol{\theta}}(\mathbf{k}^{\prime\prime})\,d\mathbf{k}^{\prime\prime}\quad\mbox{where}\quad D_{K}(\mathbf{k})=\frac{1}{2\pi}\left(\frac{\Delta x\Delta y}{MN}\right)^{\frac{1}{2}}\sum_{\mathbf{x}}w(\mathbf{x})e^{-i\mathbf{k}\cdot\mathbf{x}}.$ (61)
@@
$\langle H(\mathbf{k})H(\mathbf{k^{\prime}})\rangle=\iint\limits_{\mathbf{k}}D_{K}(\mathbf{k}-\mathbf{k}^{\prime\prime})D_{K}(\mathbf{k}^{\prime}+\mathbf{k}^{\prime\prime})\hskip 1.00006pt{\mathcal{S}}_{\boldsymbol{\theta}}(\mathbf{k}^{\prime\prime})\,d\mathbf{k}^{\prime\prime}.$ (64)
This procedure requires one Fast Fourier Transform per diagonal increasingly offset from the main diagonal in $(\mathbf{k},\mathbf{k^{\prime}})$ space, further multiplied with $\bar{m}_{\theta}(\mathbf{k})/\bar{\mathcal{S}}_{\boldsymbol{\theta}}(\mathbf{k})$ and $\bar{m}_{\theta}(\mathbf{k^{\prime}})/\bar{\mathcal{S}}_{\boldsymbol{\theta}}(\mathbf{k^{\prime}})$ and normalized into eq. (56). Like the procedure in eqs (58)–(60), it is exact [[, as opposed to Riemann discretization of eqs (61) and (64), as originally proposed by]]Guillaumin+2022.
Figure 6: Sampling and windowing effects on the covariance between periodograms: blurring and correlation illustrated for a Matérn model with parameters $\boldsymbol{\theta}_{0}=({\sigma^{2}},\nu,\rho)=(1,0.5,2)$ on a grid of $M=N=9$ pixels of size $\Delta x=\Delta y=1$ in arbitrary units. (Top row) The spatial covariance ${\mathcal{C}}_{\boldsymbol{\theta}}(\mathbf{x},\mathbf{x^{\prime}})$ on the grid that gives rise to the periodogram covariance $\mbox{cov}\big\{|H(\mathbf{k})|^{2},|H(\mathbf{k^{\prime}})|^{2}\big\}$, which is the sum of two (pseudo-)covariance terms, see eq. (57). The off-diagonal terms express the correlation between periodograms at different frequencies; the diagonal terms embody their blurring. (Bottom row) The squared blurred spectral density $\bar{{\mathcal{S}}}_{\boldsymbol{\theta}}^{2}\hskip-1.00006pt(\mathbf{k})$ is close to the variance of the periodograms $\mbox{var}\big\{|H(\mathbf{k})|^{2}\}$, as shown in the last three panels, which are identical to the diagonals of the panels above them, wrapped into a two-dimensional spectral grid.
Assembling all the various pieces, we now have three different methods to obtain the desired estimation variance in eq. (56): likelihood gradient sampling, direct calculation via the DFT matrix, and per-diagonal Fourier transformation of the spatial covariance. The first is fast and has shown good behavior in all of our tests, see Sec. 4.5. The second and third produce identical results, with the matrix implementation being memory intensive, and the per-diagonal sequential calculation still computationally heavy as it calls on $MN$ FFTs, each of which requiring $MN\log(MN)$ operations. As shown in Fig. 6, not all interaction terms are large, which has inspired randomized approaches to capture their contributions [60], e.g., via Markov-Chain Monte Carlo [45, 35] or importance sampling. Managing the interplay of the stochastic process being sampled and the characteristics of the sampling strategy is amenable to formal treatment via the significant correlation contribution as developed by [26, 28].
@@ 4.5 Synthetic numerical examples
Fig. 7 is a representative illustration of the behavior of the estimator $\hat{\theta}$ of eqs (46)–(48), in each of the Matérn parameters $\sigma^{2}$}, $\nu$, and $\rho$. We conducted five hundred inversions (484 of which successfully yielded estimates) on independently generated simulations (via circulant embedding of the spatial covariance), and studied the empirical distribution of the estimates, compared to the theoretical expression of eq. (55) using the exact method of eqs (58)–(60). Invariably, the estimates are nearly unbiased, and nearly universally Gaussian distributed, as can be seen from the histograms and the quantile-quantile plots. The top row shows the smoothly estimated [5] standardized probability density functions (shaded) of the values recovered in this experiment of sample size $101\times 111$. The abscissas were truncated to within $\pm$3 of the empirical standard deviation about the origin at the mean estimate; the percentage of the values captured by this truncation is listed in the top left of each graph. The thick gray line summarizes the empirical distributions as a Gaussian with the observed mean and variance in this set of simulations. The thick black line uses the corresponding theoretical values calculated from eq. (55). The ratio of the empirical to the theoretical standard deviation is shown listed as $\mathsf{s}/\sigma$ for each of the parameters. The dotted black line marks the true parameter values. The bottom row shows the quantile-quantile plots of the theoretical (abscissa, horizontal) versus the empirical (ordinate, vertical) distributions. The averages of the recovered values ${\sigma^{2}},\nu$, and $\rho$ are listed at the top of the second row of graphs. The true parameter values $\sigma^{2}_{0},\nu_{0}$, and $\rho_{0}$ are listed at the bottom. We attribute the slight asymmetry of the histograms and the small bends at the edges of the quantile-quantile plot to small-sample effects on minimum/maximum order statistics, and to all Matérn parameters being necessarily positive.
@@
[…]
Figure 8: An example of a likelihood (eq. 48) surface for simulated data from the experiment reported in Fig. 7, shown for one estimate very close to the truth. Every two-dimensional panel is a slice at the value of the third estimated parameter. We draw and shade ten contours in equal increments, from the likelihood at the estimate in the center, to the likelihood of the parameter pair two observed standard deviations out, in the positive quadrant of the ensemble.
Figure 9: Maximum-likelihood estimation statistics for the ensemble of 484 simulation and recovery experiments reported in Fig. 7. The mean estimate is highlighted by the gray triangle and two observed standard deviations marked gray by lines. The heavy gray ellipse is the 68% confidence region based on the ensemble. The heavy black ellipse is the predicted 68% confidence region based on the covariance predicted from eq. (55), shown and evaluated at the mean estimate, which is a close match for the set.
Figure 10: Comparison of the covariance predicted via eq. (55) and the covariance observed on the set of experiments reported in Figs 7 and 9. Shown are the relevant correlation matrices between the estimators for the three Matérn parameters $\sigma^{2}$}, $\nu$, and $\rho$, highlighting the relatively strong trade-off between $\nu$ and $\rho$.
Fig. 9 summarizes the same experiment as cross plots of the recovered parameters with their empirical summaries [3] as the ellipsoidal 68% confidence region (thick gray line), and showing the equivalent error ellipses from the theoretically predicted covariances (thick black line), which match the observations closely, albeit not perfectly. Every panel shows all pairs of parameter estimates (gray circles), with the mean marked as a gray triangle, and a cross indicating twice the standard deviation for each of the estimates. In the lower left of each graph we display the percentage of the estimates that falls within the error ellipse calculated for the observed covariance.
Fig. 10 is an alternative graphical rendition of the theoretically predicted and empirically observed parameter-estimate correlation matrix, that is, their covariance matrix normalized to unity along the diagonal. Parameter trade-offs are inherent in the Matérn description (see, for example, eq. 14). The theoretically calculated correlations are $\{{\sigma^{2}},\nu\}=-0.3714$, $\{{\sigma^{2}},\rho\}=0.8906$, and $\{\nu,\rho\}=-0.6701$, which are good predictors for those observed, as the empirical values are $\{{\sigma^{2}},\nu\}=-0.4195$, $\{{\sigma^{2}},\rho\}=0.8772$, and $\{\nu,\rho\}=-0.7259$, respectively.
@@ 4.6 Analysis of residuals
The terms ${\mathcal{S}}_{\boldsymbol{\theta}}^{-1}\hskip-1.00006pt(k)\hskip 1.00006pt|H(\mathbf{k})|^{2}$ and $\bar{\mathcal{S}}_{\boldsymbol{\theta}}^{-1}\hskip-1.00006pt(\mathbf{k})\hskip 1.00006pt|H(\mathbf{k})|^{2}$ that have appeared above in the expressions for the likelihoods (39), (48), and their derivatives (40) and (52) contain the ratio of the observed periodogram of the data to the (blurred) spectral density predicted under the model. Since the spectral density is a ‘scale’ parameter (as opposed to a ‘location’) this ratio has the usual interpretation as a measure of misfit. We have previously noted that if the Gaussian model fits, the expectation of this quadratic is $\langle\bar{\mathcal{S}}_{\boldsymbol{\theta}}^{-1}\hskip-1.00006pt(\mathbf{k})\hskip 1.00006pt|H(\mathbf{k})|^{2}\rangle=1$. Here we follow [58] to maintain that twice this quantity, a quadratic form of Gaussian random variables, given that ${\mathcal{H}}(\mathbf{x})$ and thus $H(\mathbf{k})$ are assumed to be Gaussian and uncorrelated, should be a chi-squared random variable with two degrees of freedom,
$X_{\boldsymbol{\theta}}(\mathbf{k})=\bar{\mathcal{S}}_{\boldsymbol{\theta}}^{-1}\hskip-1.00006pt(\mathbf{k})\hskip 1.00006pt|H(\mathbf{k})|^{2}\sim\chi^{2}_{2}/2.$ (65)
@@
It is imprudent to ignore and impossible to overstate the importance of such a hypothesis test. Apart from serious numerical instability and potential run-away effects, possibly caused by improper initialization of or unrealistic constraints on the optimization procedure, maximum–likelihood inversion will always return the parameter set with maximum likelihood. But whether the most likely model is, in fact, any good, then remains to be ascertained. Establishing whether eq. (65) in fact holds can be carried out visually, by inspection of the overlay of the histograms of $2X_{\boldsymbol{\theta}}$ across all wave vectors with the probability density function $\chi^{2}_{2}$, and by making ‘quantile-quantile’ plots of the ranked values of $2X_{\boldsymbol{\theta}}$ versus the inverse cumulative density function of $\chi^{2}_{2}$ evaluated at their corresponding fractional ranks. Moreover, the two-dimensional map of $X_{\boldsymbol{\theta}}(\mathbf{k})$ should show no residual structure, and will contain information on possible wavenumber ranges or specific directions (for example, in the presence of anisotropy) in which the data might be over- or under-fit. All three such representations of model quality (histograms, quantile-quantile ‘Q-Q’ plots, and two-dimensional wavenumber maps of the residuals) must be thoroughly scrutinized.
Beyond visual inspection, it is desirable to design a formal test for when the hypothesis of isotropic Matérn behavior needs to be abandoned, and the veracity of the parameters recovered by likelihood maximization called into question, regardless of how narrow their uncertainty intervals based on eq. (55) may be. Failing the test could be due, for example, to non-Gaussianity, or to the presence of patterns or preferred directions indicating that the data should rather be interpreted under anisotropic [[, e.g.,]]Goff+89b,Herzfeld+99,Olhede2008,Olhede+2014b extensions of the model. We save developing alternative hypotheses for future work.
@@
The smoothness and boundedness of the spectrum $\bar{\mathcal{S}}_{\boldsymbol{\theta}}(\mathbf{k})$, the presumed independence of $X_{\boldsymbol{\theta}}(\mathbf{k})$ between wavenumbers, and the central limit theorem should help the variable $s^{2}_{X}$ to converge to a normal variate. The central moments of the $p$th power of chi-squared variables with $m$ degrees of freedom [13] satisfy $\big\langle\!\left[\chi^{2}_{m}\right]^{p}\!\big\rangle=2^{p}\Gamma(p+m/2)/\Gamma(m/2)$, from which we obtain
@@
The case $p=1$ discussed by [58] in reducing eq. (42) to eq. (43), is easily verified. Evaluating eq. (67) for the case $p=2$ then yields the expectation of eq. (66), our test statistic,
@@
Hence we deduce that our chosen metric converges ‘in law’ [17] to a variable distributed as:
@@
[…]
@@
Fig. 12 validates the distribution of the test statistic through a series of simulation experiments with the same parameters as those used for Fig. 11. Reporting on the ‘test of the test’, specifically, whether eq. (71) holds, the two panels illustrate the distribution of $s^{2}_{X}$ across 175 realizations. Sample mean and variance of the test statistic are labeled above the first panel, which displays the histogram of $s^{2}_{X}$ over the simulations. The second panel shows the linearity of the Q-Q plots. Our conclusion is that using $s^{2}_{X}$ as a statistic results in a useful and sensitive test on the appropriateness of the Matérn model, whatever its parameters, and irrespectively of their confidence intervals.